40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 15 00:40:19 2019, gautam, Update, SUS, ETMY diagnosis plan ETMY_sensorSpectra.pdf
    Reply  Wed May 15 10:57:57 2019, gautam, Update, SUS, EY chamber opened 
       Reply  Wed May 15 17:46:24 2019, gautam, Update, SUS, ETMY inspection firstContactFiber.JPGETMY_sensorSpectra.pdf
          Reply  Wed May 15 19:36:29 2019, Koji, Update, SUS, ETMY instepction 
             Reply  Thu May 16 13:07:14 2019, gautam, Update, SUS, First contact residue removal etmy_sensors.pdfetmy_BRmode.pdf
                Reply  Thu May 16 23:31:55 2019, gautam, Update, SUS, ETMY suspension characterization 6x
                   Reply  Fri May 17 17:01:08 2019, gautam, Update, SUS, ETMY suspension characterization etmy_oplevs.pdf
          Reply  Mon May 20 11:33:46 2019, gautam, Update, SUS, ITMY inspection 
             Reply  Mon May 20 17:12:57 2019, gautam, Update, SUS, ETMY LL adjustment 
                Reply  Mon May 20 22:06:07 2019, gautam, Update, SUS, ITMY also kicked 
                   Reply  Tue May 21 00:15:21 2019, gautam, Update, SUS, Main objectives of vent achieved (?) etmy_oplevs_20190520.pdfETMY_cplxTF.pdfETMY_diagComp.pdfScreen_Shot_2019-05-21_at_12.37.08_AM.png
                      Reply  Tue May 21 21:33:27 2019, gautam, Update, SUS, ETMY HR face cleaned 
                         Reply  Wed May 22 11:53:50 2019, gautam, Update, SUS, ETMY EQ stops backed out ETMY_eigenmodes.pdf
Message ID: 14611     Entry time: Wed May 15 17:46:24 2019     In reply to: 14610     Reply to this: 14612   14623
Author: gautam 
Type: Update 
Category: SUS 
Subject: ETMY inspection 

I setup the usual mini-cleanroom setup around the ETMY chamber. Then I carried out the investigative plan outlined here.

Main finding: I saw a fiber of what looks like first contact on the bottom left (as viewed from HR side) of ETMY, connecting the optic to the cage. See Attachment #1. I don't know that this can explain the problem with the missing eigenmode, it's not a hard constraint.  Seems like something that should be addressed in any case. How do we want to remove this? Just use a tweezer and pull it off, or apply a larger FC patch and then pull it off? I'm pretty sure it's first contact and not a piece of PEEK mesh because I can see it is adhered to the HR side of the optic, but couldn't capture that detail in a photo.

There weren't any obvious problem with the magnet positioning inside the OSEM, or the suspension wire. All the EQ stop tips were >3mm away from the optic.

I also backed out the bottom EQ stops on the far (south side) of the optic by ~2 full turns of the screw. Taking another free-swinging dataset now to see if anything has changed. I will upload all the photos I took, with annotations, to the gPhotos later today eve. Light doors back on at ~1730.

Update 10pm: the photos have been uploaded. I've added a "description" to each photo which should convey the message of that particualr shot, it shows up in my browser on the bottom left of the photo but can also be accessed by clicking the "info" icon. Please have a look and comment if something sticks out as odd / requires correction.

Update 1045pm: I looked at the freeswinging data from earlier today. Still only 3 peaks around 1 Hz.

The following optics were kicked:
ETMY
Wed May 15 17:45:51 PDT 2019
1242002769
Attachment 1: firstContactFiber.JPG  834 kB  Uploaded Wed May 15 20:02:20 2019  | Hide | Hide all
firstContactFiber.JPG
Attachment 2: ETMY_sensorSpectra.pdf  142 kB  Uploaded Wed May 15 23:39:57 2019  | Hide | Hide all
ETMY_sensorSpectra.pdf
ELOG V3.1.3-